Father is 3 times as old as his daughters age.In 12 years father will be 1 year less than twice his daugthers age.What is their age?
Answer by stanbon(75887)

Father is 3 times as old as his daughters age.In 12 years father will be 1 year less than twice his daugthers age.What is their age?
--------------
Equations:
f = 3d
f+12 = 2(d+12)-1
-----
Substitute for "f" and solve for "d":
3d+12 = 2d+23
d = 11 (daughter's age now)
----
Solve for "f":
f + 3d
f = 33 (father's age now)
=================
Cheers,
Stan H.
=========